SELF- AND H_{2}-BROADENING AND SHIFT COEFFICIENTS IN THE 2 ← 0 BAND OF^{12}C^{16}O: REVISITED

V. MALATHY DEVI, D. CHRIS BENNER, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795; A. PREDOI-CROSS, Department of Physics, The University of Lethbridge, Lethbridge, AB T1K 3M4, Canada; \underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Hampton, VA 23681-2199; and A. W. MANTZ, Department of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320.

Room temperature values for self-broadened and hydrogen-broadened Lorentz half width coefficients, and self and hydrogen pressure-induced shift coefficients have been measured for transitions with rotational quantum number index m ranging from −24 to +24 in the 2 ← 0 band of ^{12}C^{16}O. The spectra were recorded with the McMath-Pierce Fourier transform spectrometer located at the National Solar Observatory on Kitt Peak. The analysis was performed using a multispectrum nonlinear least squares technique^{a} modified to constrain the Lorentz widths in the P and R branches to be identical for the same |m| value. We have compared the present results with our previous measurements made with the same spectrometer^{b} and with other measurements published recently.

aD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

bV. Malathy Devi, D. Chris Benner, M. A. H. Smith, C. P. Rinsland and A. W. Mantz, JQSRT \underline{75}, 455-471 (2002).